Circular No. 8672

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2005 W16, W17, X2–X9, Y3–Y6 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8657). According to K. Battams, C/2005 W16 and C/2005 X5 were small and stellar in appearance in C3 images, the former reaching mag 7.3 at 10.1R_{⊙} on Nov. 28.471 UT, and the latter reaching mag 6.4 at 9.3R_{⊙} on Dec. 8.846; in C2 images, C/2005 W16 showed a faint, thin tail that reached a length of ∼ 9′ on Nov. 28.854 at 5.7R_{⊙}, while C/2005 X5 was fairly diffuse with a very faint tail that reached a length of ∼ 6′ on Dec. 9.083 at 6.7R_{⊙}. C/2005 W17, C/2005 X2, and C/2005 X6 were tiny and too faint for photometry; C/2005 W17 was diffuse and the latter two were stellar in appearance. C/2005 X3 was small and diffuse, reaching mag 7.7 at 6.5R_{⊙} on Dec. 4.600. C/2005 X4 was small, faint, and diffuse, reaching mag ∼ 8. C/2005 X7 was tiny and stellar in appearance, yet persistent (despite reaching mag ∼ 8), curiously surviving as long as did C/2005 X5. C/2005 X8 and C/2005 Y5 were extremely faint — the former described as tiny, and the latter as small and diffuse; C/2005 Y6 was a near-identical twin to C/2005 Y5, but marginally brighter. C/2005 X9 was small and diffuse, reaching mag 7.7 at 7.3R_{⊙} on Dec. 12.917. C/2005 Y3 was small and a little too diffuse for good photometry in C3 images; in C2 images, it was also diffuse, reaching mag 6.6 at 7.4R_{⊙} on Dec. 20.379. C/2005 Y4 was very faint and diffuse.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 W16 Nov. 28.113 16^{h}09.^{m}4 −25^{o}06′ C3/2 RM 2006-C59
C/2005 W17  29.704 16 23.1 −23 19 C2  HS 2006-C59
C/2005 X2  Dec. 3.442 16 40.5 −23 55 C2  HS 2006-C59
C/2005 X3  4.583 16 45.6 −24 09 C2  HS 2006-C59
C/2005 X4  6.567 16 54.5 −24 21 C2  KC 2006-C60
C/2005 X5  8.321 16 59.0 −26 59 C3/2 JR 2006-C60
C/2005 X6  8.392 17 03.0 −24 37 C2  JR 2006-C60
C/2005 X7  9.075 17 07.9 −24 46 C2  HS 2006-C60
C/2005 X8  11.908 17 19.5 −24 58 C2  QY 2006-C61
C/2005 X9  12.908 17 23.9 −25 03 C2  RM 2006-C61
C/2005 Y3  19.988 17 57.4 −26 25 C3/2 QY 2006-C61
C/2005 Y4  20.938 18 01.5 −25 11 C2  SH 2006-C61
C/2005 Y5  22.354 18 07.5 −25 13 C2  KB 2006-C61
C/2005 Y6  22.454 18 07.9 −25 10 C2  KB 2006-C61
2006 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green